# SLR: Learning Quadruped Locomotion without Privileged Information

### v1

@@ Abstract
Traditional reinforcement learning control for quadruped robots often relies on privileged information, demanding meticulous selection and precise estimation, thereby imposing constraints on the development process. This work proposes a Self-learning Latent Representation (SLR) method, which achieves high-performance control policy learning without the need for privileged information. To enhance the credibility of our proposed method’s evaluation, SLR is compared with open-source code repositories of state-of-the-art algorithms, retaining the original authors’ configuration parameters. Across four repositories, SLR consistently outperforms the reference results. Ultimately, the trained policy and encoder empower the quadruped robot to navigate steps, climb stairs, ascend rocks, and traverse various challenging terrains. Robot experiment videos are at https://11chens.github.io/SLR/
@@
Figure 1: We present a framework designed to train a robust quadruped locomotion policy, eliminating the necessity for privileged information. The robot adeptly maneuvers through diverse terrains in simulation and exhibits comparable performance in corresponding real-world environments, showcasing a remarkable level of proficiency.
@@ 1 Introduction
Most RL applications in quadruped robots rely on privileged learning methods [19]. In real-world scenarios, a robot’s interaction with its environment is modeled as a Partially Observable Markov Decision Process (POMDP). Solely relying on proprioceptive sensor measurements, a robot cannot fully perceive external environmental information, limiting its decision-making capabilities. Consequently, many studies leverage the “observability” advantages of simulation platforms. During training, various physical parameters (such as friction coefficients [10, 20, 21], restitution coefficients [9, 21], and scan-dots of the terrain [14, 22]) are artificially added as privileged information to help the robot understand both itself and the external environment.
In contrast, human cognition does not require explicit knowledge of physical parameters to traverse various terrains. Similarly, for robots using neural networks as their strategy, adding specific physical parameters may not yield the anticipated results. This discrepancy arises because we are not solving dynamic equations based on robot models but incorporating them as part of the neural network’s input. Hence, these crafted privileged pieces of information may not necessarily provide comprehensibility or interpretability for neural network-based agents.
Therefore, instead of relying on manually chosen physical parameters to construct privileged information, this work explores whether it is possible for robots to learn a latent representation of environmental states by themselves? We believe that sufficiently intelligent robots should possess this capability, and representations learned through self-learning are more suitable for robots than those specified by humans.
To this end, a Self-learning Latent Representation (SLR) algorithm is proposed that generates latent representations guided by the Markov process of reinforcement learning without using any privileged information. Through self-learning, the robot can understand the latent features of the environment, demonstrating generalized locomotion capabilities across various terrains,as shown in Figure 1.
Our results indicate that, without relying on any privileged information, the SLR algorithm outperforms traditional privileged learning methods. Moreover, the learned latent representations are highly consistent with the actual terrain conditions across various terrains. To evaluate the proposed algorithm, the SLR algorithm is implemented in the open-source code repositories of previous researchers and compared in the same environments used by the authors. The SLR results demonstrated that this approach achieves state-of-the-art performance both in simulation and real-world deployments.
@@ 3.1 Problem Formulation
Figure 2: Illustration of SRL training framework. All dashed lines represent the network updating process. The solid pink line indicates the encoder updates through backpropagation from the Critic network, the transition model, and random sampling. The remaining solid lines represent the network’s forward inference process.
@@ 3.2 Framework Overview
The proposed training framework fully leverages the Markov Decision Process (MDP), guiding the latent’s self-learning based on state transitions, state distinctions, and cumulative rewards, without relying on manually set privileged information constraints. All networks used in this framework are multi-layer perception (MLP). The training framework is shown in Figure 2.
Encoder: In this architecture, the encoder’s input consists of the observation history $o^{H}_{t}$, which is composed of proprioceptive information $o_{t}$ from the previous 10 time steps. The encoder outputs a latent representation $z_{t}$ of the observation history:
@@
State Transition Model: The state transition model simulates the real state transitions of the environment $p\left(s_{t+1}\mid s_{t},a_{t}\right)$ and shares the same network structure as the encoder. Its input is the latent representation $z_{t}$ and the action $a_{t}$, and it outputs the next time step’s latent estimation $\tilde{z}_{t+1}$:
@@
where $m$ is a margin that is enforced between $\tilde{z}_{t+1}$ and $z_{t+n}$ pairs.
This updating strategy empowers the encoder to comprehend the dynamics of environmental state transitions and extract environmental attributes by assimilating state-action pairs derived from the MDP rollout.
@@ 4 Experiments
4.1 Ablation Study for the Privileged Learning Encoder
To evaluate the proposed algorithm against traditional privileged learning algorithms, an ablation study was conducted on Implicit and Explicit privileged learning methods, as well as the SLR method. Commonly used privileged information $e_{t}\in\mathbb{R}^{10}$ were chosen, including parameters such as friction, restitution, foot height, and foot contact. During training, the former two parameters were utilized for domain randomizations to enhance model generalization, while the latter two served as external robot perceptions aiding decision-making. The above two privileged learning methods are as follows:
Implicit: We adapted the teacher algorithm from [9] to encode privileged information into latent $l_{t}=h(e_{t})$, which was integrated into the policy. An adaptation module then estimated the privileged latent from observation history $\tilde{l}_{t}=\phi(o^{H}_{t})$ and aligned it with the real privileged latent generated by the encoder.
Explicit: The 10-dimensional privileged information ${e}_{t}$ was directly estimated as $\tilde{e}_{t}=\phi(o^{H}_{t})$ and then integrated into the policy.
To ensure the fairness of quantitative comparisons, the configurations used in the experiments, including training hyperparameters, reward function, and other settings, are all based on the default settings in the code repository [27]. Training was conducted for 5000 iterations under multiple terrains by default, and the results are presented in Figure 3.
Figure 3: Ablation study training curves, curves are averaged over 3 seeds. The shaded area represents the standard deviation across seeds.
Figure 3 demonstrates the significant superiority of the proposed algorithm over traditional privileged learning methods. This improvement is attributed to the self-learning encoding mechanism, enabling the robot to discern between various terrains. While certain privileged information, such as friction and restitution used for domain randomization, remains consistent across different terrains, others like foot height and foot contact offer limited terrain differentiation.
@@ 4.2 Latent Representation Analysis
Identifying and distinguishing various terrain types is essential for robotic systems [28]. To assess the efficacy of the self-learned latent representation in this regard, a simulation-based latent representation analysis is conducted. A test environment is designed composed of four sequential terrains: an upward slope, descending stairs, flat ground, and ascending stairs. The trained robot is tasked with navigating these terrains sequentially (Figure 4). During traversal, the latent representation output is recorded at each step. Subsequently, t-distributed stochastic neighbor embedding (t-SNE) analysis is applied to project the complex latent representations of the entire trajectory into a two-dimensional space for analysis. The latent representations from Implicit and SRL are analyzed separately, as depicted in Figure 5.
@@
Figure 4: t-SNE test terrain.
@@ 5.1 Compared Methods
In this study, these issues are addressed by directly implementing the SLR algorithm in the opensource code repositories of previous state-of-the-art (SOTA) works [9, 11, 13, 27]. Only the algorithm framework is modified while keeping other variables consistent with the original settings. This approach ensures a fair comparison between our method and the publicly available SOTA algorithms. The algorithms which are evaluated as given below:
  • MoB: Explicit estimation, trained on flat ground with the Unitree Go1 robot.
  • RLvRL: Implicit estimation, trained on flat ground with the Cheetah robot.
  • ROA: Implicit estimation, trained on custom fractal noise environment using a Unitree Go1 robot equipped with a manipulator.
  • Baseline: No encoder, no privileged information, trained in multi-terrain environments with the Unitree A1 robot.
@@ 5.2 Simulation
All four code repositories implemented RL based on the PPO algorithm, with training conducted on the IsaacGym platform [27, 29], utilizing three different random seeds. Training curves are illustrated in Figure 6. Evaluation metrics include linear velocity tracking error (LVTE) and angular velocity tracking error (AVTE), assessed using Mean Square Error (MSE). Target linear velocity ranges from -1.0 to 1.0 m/s, and target angular velocity ranges from -1.0 to 1.0 rad/s. Evaluation results are summarized in Table 1.
Figure 6: Training curves from various repositories. The SRL’s mean reward surpasses those of Baseline, RLvRL, and ROA, and is marginally higher than MoB. The shaded area represents the standard deviation across seeds.
Table 1: Velocity tracking errors in various environments. The SRL algorithm demonstrates superior velocity tracking capabilities compared to the authors’ algorithm in most cases. Top performances are highlighted in bold.
Metrics Method Code Repository
MoB RLvRL ROA Baseline
LVTE Author 0.078 0.051 0.150 0.041
Ours 0.115 0.080 0.136 0.033
AVTE Author 0.130 0.229 1.100 0.053
Ours 0.041 0.125 1.096 0.045
Based on the evaluation results of the four code repositories, the proposed SLR algorithm generally achieves higher mean rewards and superior velocity tracking capabilities compared to the original implementations. This indicates that this approach can develop more effective policies without relying on privileged information. Additionally, the experiments with the MoB and ROA repositories demonstrate that the SLR algorithm is robust and generalizable, capable of handling complex tasks involving multiple commands and manipulators. Consequently, it can be inferred that enabling robots to autonomously learn the latent representation yields greater benefits than manually selecting privileged information.
@@ 5.3 Deploy in Real-World
The trained policies are deployed on the Unitree Go2 robot in real-world as depicted in Figure 7. Performance evaluation of the policy was conducted across various indoor and outdoor terrains, and comparative analyses were performed against above code repositories and Unitree built-in MPC control method. Each environment was tested ten times. As presented in Table 2, the results demonstrate the superior efficacy of our policy in real-world scenarios.
@@ 6 Discussion and Limitations
In this study, we introduce a quadruped locomotion algorithm that operates without relying on privileged information. Unlike previous open-source repositories, our proposed algorithm surpasses the author’s algorithm in performance retaining original authors’ configuration parameters. Through real-world deployment, our robot demonstrates the capability to navigate through diverse and challenging terrains.
@@
Figure 7: Indoor and outdoor experiment settings. The left column (a) showcases indoor scenarios, with the top row depicting the robot navigating a step obstacle and the bottom row featuring the robot on a deformable surface. The right column (b) illustrates outdoor scenarios, depicting the robot moving through a bush, on a cobbled road, and climbing a curb in the top row, while the bottom row showcases the robot encountering a slope, stairs, and a rock.
Table 2: Comparison of different deployment strategies across various terrains. Each strategy was evaluated ten times per terrain, with the values in the table indicating the number of successful trials. Top performances are highlighted in bold.
Scenarios Terrain Ours MoB RLvRL Baseline MPC
Indoor Step 10 4 2 1 10
Deformable Slope 10 3 1 0 8
Outdoor Bush 10 9 7 5 10
Cobbled Road 10 10 4 3 9
Curb 10 2 1 0 4
Earthen Slope 10 4 2 0 6
Stairs 10 0 0 0 2
Rock 7 0 0 0 0

### v2

@@ Abstract
The recent mainstream reinforcement learning control for quadruped robots often relies on privileged information, demanding meticulous selection and precise estimation, thereby imposing constraints on the development process. This work proposes a Self-learning Latent Representation (SLR) method, which achieves high-performance control policy learning without the need for privileged information. To enhance the credibility of the proposed method’s evaluation, SLR was directly compared with state-of-the-art algorithms using their open-source code repositories and original configuration parameters. Remarkably, SLR surpasses the performance of previous methods using only limited proprioceptive data, demonstrating significant potential for future applications. Ultimately, the trained policy and encoder empower the quadruped robot to traverse various challenging terrains. Videos of our results can be found on our website: https://11chens.github.io/SLR/
@@
Figure 1: We propose a framework for training a robust quadruped locomotion policy without relying on privileged information. The robot effectively navigates challenging terrains, showcasing adaptive locomotion skills acquired through self-learning.
@@ 1 Introduction
The recent mainstream RL applications in quadruped robots rely on privileged learning methods [19]. In real-world scenarios, a robot’s interaction with its environment is modeled as a Partially Observable Markov Decision Process (POMDP). Solely relying on proprioceptive sensor measurements, a robot cannot fully perceive external environmental information, limiting its decision-making capabilities. Consequently, many studies leverage the “observability” advantages of simulation platforms. During training, various physical parameters (such as friction coefficients [10, 20, 21], restitution coefficients [9, 21], and scan-dots of the terrain [14, 22]) are artificially added as privileged information to help the robot understand both itself and the external environment.
Unlike human cognition, which navigates terrains without explicit knowledge of physical parameters, neural network-based robots may not benefit from adding such parameters. This is because the parameters are not used to solve dynamic equations but are simply input into the neural network. As a result, these privileged information pieces may not enhance the interpretability or effectiveness of neural network-based agents.
Therefore, instead of relying on manually chosen physical parameters to construct privileged information, this work explores whether it is possible for robots to learn a latent representation of environmental states by themselves? It is proposed that intelligent robots should be capable of self-learning latent representations, which are inherently more suitable than those predefined by humans.
To achieve this, we propose the Self-learning Latent Representation (SLR) algorithm, which generates latent representations guided by a RL Markov process without privileged information. This self-learning approach enables the robot to grasp latent environmental features and exhibit generalized locomotion capabilities across challenging terrain, as illustrated in Figure 1.
Our results show that the SLR algorithm, which operates without privileged information, outperforms traditional privileged learning methods. It consistently aligns with actual terrain conditions across various terrains. Implemented in existing open-source repositories and evaluated in the same environments as previous studies, the SLR algorithm demonstrates state-of-the-art (SOTA) performance in both simulations and real-world applications. This suggests that the self-learning approach has considerable potential for future expansion.
@@ 3.1 Problem Formulation
Figure 2: Illustration of SLR training framework. All dashed lines represent the network updating process. The translucent fuchsia lines indicate the encoder updates through backpropagation from the Critic network, the Transition model, and random sampling. The remaining solid lines represent the network’s forward inference process.
@@ 3.2 Framework Overview
The proposed training framework utilizes the Markov Decision Process (MDP) to guide self-learning of the latent representation through state transitions, distinctions, and cumulative rewards, without relying on manually defined privileged information. All networks in this framework are multi-layer perceptrons (MLPs). Table A3 in the Appendix details the MLP structures, and Table A4 lists the training hyperparameters. The training framework is illustrated in Figure 2.
Encoder: In this architecture, the encoder’s input consists of the observation history $o^{H}_{t}$, which is composed of proprioceptive information $o_{t}$ from the previous 10 time steps. The encoder outputs a 20-dimensional latent representation $z_{t}$ of the observation history:
@@
Transition Model: The Transition model simulates the real state transitions of the environment $p\left(s_{t+1}\mid s_{t},a_{t}\right)$. Its input is the latent representation $z_{t}$ and the action $a_{t}$, and it outputs the next time step’s latent estimation $\tilde{z}_{t+1}$:
@@
where $m$ is a margin that is enforced between $\tilde{z}_{t+1}$ and $z_{t+n}$ pairs, set to 1.0 in the experiment.
This updating strategy empowers the encoder to comprehend the dynamics of environmental state transitions and extract environmental attributes by assimilating state-action pairs derived from the MDP rollout. Finally, the triplet loss $\mathcal{L}_{\text{trip}}$ is multiplied by a triplet coefficient and added to the PPO loss $\mathcal{L}_{\text{ppo}}$ [24] for network updating. Details of simulation and training are in Appendix 1.
@@ 4 Experiments
4.1 Ablation Study for Latent Representations
To evaluate the proposed algorithm against traditional privileged learning algorithms, an ablation study was conducted. In this setup, commonly used privileged information $e_{t}\in\mathbb{R}^{10}$ was chosen, including parameters such as friction, restitution, foot height, and foot contact. The methods using different types of latent representations are as follows:
  1. 1)
    Implicit: The policy input includes the implicit privileged latent $l_{t}$, which is encoded from privileged information $e_{t}$ by adopting the teacher policy approach in [9, 10].
  2. 2)
    Explicit: The policy input includes the explicit privileged information $\tilde{e}_{t}=\phi(o^{H}_{t})$.
  3. 3)
    SLR w/ explicit: The SLR policy input is the concatenation of the self-learning latent $z_{t}$ and explicit privileged latent $\tilde{e}_{t}$.
  4. 4)
    SLR w/ implicit: The SLR policy input is the concatenation of the self-learning latent $z_{t}$ and implicit privileged latent $l_{t}$.
  5. 5)
    SLR w/o latent: The SLR policy input does not include any latent representations and does not utilize an encoder.
To ensure the fairness of quantitative comparisons, the configurations used in the experiments are all based on the default settings in the code repository [27]. Training was conducted for 6000 iterations under multiple terrains by default, and the results are presented in Figure 3. And we can draw the following conclusions:
1. The SLR algorithm outperforms traditional explicit and implicit privileged learning methods, suggesting that unified privileged information struggles to effectively represent diverse terrains.
2. Incorporating additional privileged latent information reduces the SLR algorithm’s performance, possibly due to increased policy input disrupting the critic’s guidance on updating the self-learning latent.
Figure 3: Ablation study training curves, curves are averaged over 3 seeds. The shaded area represents the standard deviation across seeds, and the curves are smoothed using Gaussian filtering.
@@ 4.2 Latent Representation Analysis
Identifying and distinguishing terrain types is crucial for robotic systems [28]. To evaluate the self-learned latent representation, we conducted a simulation-based analysis. The test environment includes four terrains: an upward slope, descending stairs, flat ground, and ascending stairs. The robot navigates these terrains sequentially (Figure 4), and the latent representation output is recorded at each step. We then apply t-distributed stochastic neighbor embedding (t-SNE) to project these complex latent representations into a two-dimensional space for analysis. The latent representations from Implicit and SLR are examined separately, as shown in Figure 5.
@@
Figure 4: t-SNE test terrains.
@@ 5.1 Compared Methods
In this study, these issues are addressed by directly implementing the SLR algorithm in the opensource code repositories of previous SOTA works [9, 11, 13, 27, 29]. Only the algorithm framework is modified while keeping other variables consistent with the original settings. This approach ensures a fair comparison between our method and the publicly available SOTA algorithms1. 1Previous evaluations in a single repository achieved noticeably better performance (see Section 4.1). To provide more conclusive evidence, we extended our experiments.The algorithms which are evaluated as given below:
  • MoB[13]: Explicit estimation, trained on flat ground with the Unitree Go1 robot.
  • RLvRL[9]: Implicit estimation, trained on flat ground with the Cheetah robot.
  • ROA[11]: Implicit estimation, trained on custom fractal noise environment using a Unitree Go1 robot equipped with a manipulator.
  • HIM[29]: Explicit and implicit estimation, trained in multi-terrain environments with the Unitree Aliengo robot.
  • Baseline[27]: No encoder, no privileged information, trained in multi-terrain environments with the Unitree A1 robot.
@@ 5.2 Simulation
All five code repositories implemented reinforcement learning based on the PPO algorithm, with training conducted on the IsaacGym platform [27, 30] using three different random seeds. The training curves are illustrated in Figure 6. Additionally, to evaluate the locomotion capabilities, experiments on velocity tracking performance were conducted on flat terrain. The evaluation metrics included Linear Velocity Tracking Error (LVTE) and Angular Velocity Tracking Error (AVTE), both measured using Mean Square Error (MSE). The evaluation results are summarized in Table 1.
Figure 6: Training curves from various repositories show that the mean reward of SLR surpasses those of Baseline, RLvRL, and ROA, is slightly higher than MoB, and is comparable to HIM. The shaded area represents the standard deviation across seeds, and the curves are smoothed using Gaussian filtering. Note: SLR achieved a higher terrain level than HIM, indicating superior performance.
Table 1: Velocity tracking errors in various environments. The SLR algorithm demonstrates superior velocity tracking capabilities compared to the previous algorithm in most cases. Top performances are highlighted in bold. Note: N/A indicates that the ROA code repository does not support rapid locomotion.
Metrics Ranges Method Code Repository
MoB RLvRL ROA HIM Baseline
LVTE [-1,1] m/s Previous $0.058^{\pm 0.011}$ $0.032^{\pm 0.003}$ $0.150^{\pm 0.001}$ $0.014^{\pm 0.001}$ $0.021^{\pm 0.005}$
Ours $\mathbf{0.029^{\pm 0.004}}$ $\mathbf{0.024^{\pm 0.002}}$ $\mathbf{0.136^{\pm 0.001}}$ $\mathbf{0.013^{\pm 0.002}}$ $\mathbf{0.011^{\pm 0.004}}$
[-2,2] m/s Previous $0.181^{\pm 0.014}$ $0.108^{\pm 0.006}$ N/A $0.019^{\pm 0.009}$ $0.074^{\pm 0.007}$
Ours $\mathbf{0.124^{\pm 0.023}}$ $\mathbf{0.106^{\pm 0.002}}$ N/A $\mathbf{0.017^{\pm 0.002}}$ $\mathbf{0.066^{\pm 0.004}}$
AVTE [-1,1] rad/s Previous $\mathbf{0.035^{\pm 0.002}}$ $0.182^{\pm 0.007}$ $1.100^{\pm 0.041}$ $\mathbf{0.013^{\pm 0.009}}$ $0.137^{\pm 0.013}$
Ours $0.049^{\pm 0.001}$ $\mathbf{0.057^{\pm 0.007}}$ $\mathbf{1.096^{\pm 0.025}}$ $0.014^{\pm 0.011}$ $\mathbf{0.077^{\pm 0.019}}$
[-2,2] rad/s Previous $0.131^{\pm 0.015}$ $0.347^{\pm 0.002}$ N/A $0.038^{\pm 0.019}$ $0.315^{\pm 0.021}$
Ours $\mathbf{0.086^{\pm 0.009}}$ $\mathbf{0.142^{\pm 0.002}}$ N/A $\mathbf{0.038^{\pm 0.023}}$ $\mathbf{0.223^{\pm 0.026}}$
Based on the evaluation results of the five code repositories, the proposed SLR algorithm generally achieves higher mean rewards and superior velocity tracking capabilities compared to the original implementations. From these results, we can draw the following conclusions:
1. The SLR algorithm achieves mean rewards comparable to those of the recently released top-performing HIM algorithm, but with a higher terrain level (see Figure 8), indicating superior performance. This demonstrates that learning from limited proprioceptive data can rival advanced privileged learning methods.
2. The SLR algorithm’s effectiveness in managing complex tasks with MoB (multiple commands) and ROA (manipulator-equipped) repositories suggests that self-learning latent representations offer more general benefits than manually selected privileged information.
3. The SLR algorithm demonstrates enhanced tracking capabilities across various velocity ranges. One reason for this is that SLR can implicitly infer velocity from proprioceptive data. Additionally, the Critic network optimizes for maximum cumulative rewards, with velocity tracking as the most significant term in the reward function. Consequently, the backpropagation in the Critic network naturally guides the latent representation to optimize tracking performance.
@@ 5.3 Deploy in Real-World
The trained policies are deployed on the Unitree Go2 robot in real-world as depicted in Figure 7. Performance evaluation of the policy was conducted across various indoor and outdoor terrains, and comparative analyses were performed against code repositories [13, 27, 29] and Unitree built-in MPC control method. Each environment was tested 10 times. As presented in Table 2, the results demonstrate the superior efficacy of our policy in real-world scenarios.
@@ 6 Discussion and Limitations
In this study, we present a quadruped locomotion algorithm that operates without privileged information. Relying solely on limited proprioceptive data, the algorithm achieves SOTA performance. Looking ahead, we anticipate that future self-learning approaches will be able to integrate privileged information and external perceptions, leading to even more impressive outcomes.
@@
Figure 7: Indoor and outdoor experiment settings. The left column (a) showcases indoor scenarios, with the top row depicting the long stairs and the bottom row featuring the step obstacle.The right column (b) illustrates outdoor scenarios, depicting the robot moving through a bush, on a cobbled road, and climbing a curb in the top row, while the bottom row showcases the robot encountering a slope, stairs, and a rock.
Table 2: Comparison of deployment methods across various terrains. Each method was tested 10 times per terrain. The “Number ” indicates the total stairs ($16\times 29$ cm) climbed by the robot from start to fall, and “Height” shows the maximum step height the robot can handle. The $\pm$ symbol represents a 95% confidence interval. A successful trial means the robot completes the terrain without intervention or failure within the allotted time. Top performances are in bold.
Scenarios Terrain Metrics Ours HIM MoB Baseline MPC
Indoor Stairs Number $\mathbf{136.4^{\pm 24.5}}$ $107.4^{\pm 19.1}$ $0.0^{\pm 0.0}$ $0.0^{\pm 0.0}$ $93.5^{\pm 32.9}$
Step Height (cm) $\mathbf{35.4^{\pm 2.2}}$ $30.7^{\pm 2.7}$ $6.4^{\pm 1.2}$ $5.3^{\pm 0.5}$ $15.6^{\pm 1.6}$
[…]
Cobbled Road $\mathbf{100}$ $\mathbf{100}$ $\mathbf{100}$ $30$ $90$
Curb $\mathbf{100}$ $\mathbf{100}$ $20$ 00 $40$
Earthen Slope $\mathbf{100}$ $\mathbf{100}$ $40$ 00 $60$
Stairs $\mathbf{100}$ $\mathbf{100}$ 00 00 $20$
Rock $\mathbf{70}$ $30$ 00 00 00